# Ergodic Theorems and Equivalence of Green’s Kernel for Random Walks in Random Environments

### v3

@@
Abstract
We study the Ergodic Properties of Random Walks in stationary ergodic environments without uniform ellipticity under a minimal assumption. There are two main components in our work. The first step is to adopt the arguments of Lawler to first prove a uniqueness principle. We use a more general definition of environments using Environment Functions. As a corollary, we can deduce an invariance principle for balanced environments under the assumptions. We also use the uniqueness principle to show that any balanced, elliptic random walk must have the same transience behaviour as the simple symmetric random walk. The second is to transfer the results we deduce in balanced environments to general ergodic environments(under some assumptions) using a control technique to derive a measure under which the local process is stationary and ergodic. As a consequence of our results, we deduce the Law of Large Numbers for the Random Walk and an Invariance Principle under our assumptions.
@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments (RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers. Roughly speaking, it says there is a deterministic limiting velocity of the walk. This has been studied in detail in $d=1$ by Solomon[10] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
Informally speaking, the problem regarding the Law of Large Numbers in RWRE is the following:
@@
$\frac{X_{n}}{n}\to v\quad\text{a.e.}$
for almost every choice of environment?
Varadhan in [11] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$. Furthermore, Noam Berger in [1] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0. These results assume that the environment is uniformly elliptic and i.i.d.. In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
We shall work under a more general definition of environments using Environment Function, which we shall introduce below. We shall establish under some minimal assumption that this chain has a stationary measure for balanced environments, adapting arguments from [7] for any environment function. This shall also provide an easy proof of Theorem 1 of [4]. This shall be derived by first establishing a uniqueness principle for concave functions on the lattice satisfying certain properties. We answer the following natural question for balanced random walks, which can be thought of for now as a random walk having equal probabilities of going forward or backward along each coordinate axis.
@@
Is every balanced, elliptic(probability of going along every axis at every point is positive) walk in $\mathbb{Z}^{d}$ recurrent in $d=1,2$ and transient in $d\geq 3$?
We answer this in the affirmative by comparing the walk with the simple symmetric random walk in the dimension of interest. Using the uniqueness, we shall derive bounds on the Green’s Kernel of Balanced Elliptic Walks. It is important to note that this section shall not deal with the notion of a Random Environment. This is one of the key contributions of this paper, as it shows that any balanced, elliptic random walk must be recurrent in $d=1,2$ and transient for $d\geq 3$, thereby establishing Pölya’s Theorem in this general setup.
Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process. This process chain is a projection at the origin of the “Environment Viewed from the Particle” Markov Chain or, as we will call it in this paper, the environment process with a suitable environment function. We consider the local process as environment process does not always possess suitable ergodic properties. This can be noted for Random Walks in Dirichlet Environments as studied by Christopher Sabot in [9].
Our definition of environments allows us to transfer our results from the balanced case to the general environment by coupling the random walk and a reflected walk. We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof.
As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments. To the best of our knowledge, these theorems are novel. Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose study the extent to which the condition of uniform ellipticity can be dropped.
@@ 2. Prerequisites
Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [7].
@@
$\mathcal{S}(d)$ $=\{\text{set of functions from }\mathbb{Z}^{d}\to S(d)\}.$
Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathcal{S}(d)$ respectively.
Take $\mathscr{E}:\mathcal{S}(d)\to\mathcal{S}(d)$ such that $\mathscr{E}(\omega)(x)\equiv\mathscr{E}(\omega(x)):S(d)\to S(d)$ is a measurable function on $(S(d),\mathcal{B}(d))$. We call this the environment function. This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$. This is essentially a mild but powerful generalisation of the environment, as we shall see.
@@
Given an environment function $\mathscr{E}$, $\omega\in\mathcal{S}(d)$, we call the pair $(\omega,\mathscr{E})$ as the environment.
Let $\mu$ be a probability measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
@@
Define the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathcal{S}(d)$ given by
@@
For $\mathscr{E}(\omega)\in\mathcal{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
@@
Fix $\omega\in\mathcal{S}(d)$ and an environment function $\mathscr{E}$. Define the Markov Chain of the Random Walk $\{X_{\mathscr{E}(\omega)}(k)\}_{k\geq 0}$ in $\mathbb{Z}^{d}$ by the transition probabilities,
@@
Define the environment process as the Markov Chain given by $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with the transition kernel given by
@@
Then we define the local process associated with any discrete time Markov Chain $\{Y_{k}\}_{k\geq 0}$ on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ as $\{\pi_{0}(Y_{k})\}_{k\geq 0}$.
@@
Therefore if ${\mathscr{E}(\omega)}\in\mathcal{S}(d)$ one has that
$\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}(G):\mathcal{S}(d)\to[0,1]$
is measurable $\forall G\in\mathcal{G}$. Hence, we can define the annealed law to be the probability measure on $(\mathcal{S}(d)\times(\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{B}(d)^{\prime}\times\mathcal{G)}$ as
@@
We make the following assumptions on the environment that is
@@
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Define the quantity $c(\mathscr{E},x)$ as follows
@@
We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$. A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{Z}^{d}$ $(r>0)$ is said to be concave if
@@ 3. Main Results and Discussions
We first state the result that allows us to effectively deal with environments beyond uniform ellipticity. We shall demonstrate this through another result later in this section.
@@
Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
Then there exists a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
@@
Let $d\geq 1$. Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function. Fix $\omega\in\mathcal{S}(d)$. Then under assumptions 2.2 and 2.3,
@@
These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [4] for $d\geq 1$, which states that
Theorem 3.3 (Theorem 1(i) of [4]).
@@
$c(x)=\prod_{i=1}^{d}[\omega(x,e_{i})]^{\frac{1}{d}}.$
@@
To prove the above results, we derive a uniqueness theorem for concave functions on $\mathbb{Z}^{d}$, which also gives the following theorem on the transience of balanced elliptic walks. The result compares the walk to the simple symmetric random walk on $\mathbb{Z}^{d}$ given by the law
$\mathbb{P}[X_{1}=x+v\Big|X_{0}=x]=\frac{1}{2d}.$
To state our result, we recall the so-called Green’s function for a random walk on $\mathbb{Z}^{d}$.
Definition 3.4.
Let $(\omega,\mathscr{E})$ be a balanced environment in $\mathbb{Z}^{d}$($d\geq 1$). Then the Green’s function $G^{\mathscr{E}(\omega)}:\mathbb{Z}^{d}\times\mathbb{Z}^{d}\to\mathbb{R}^{+}\cup\{0\}$ is given by
$G^{\mathscr{E}(\omega)}(x,y)=\mathbb{E}_{x}[\sum_{j=0}^{\infty}\boldsymbol{1}_{X_{\mathscr{E}(\omega)}(j)=y}],$
where the expectation is taken conditional on $X_{\mathscr{E}(\omega)}(0)=x$. For the simple symmetric random walk, we shall denote its Green’s function as just $G$.
Theorem 3.5.
Let $d\geq 1$. Let $(\omega,\mathscr{E})$ be any balanced elliptic environment. Then
$\frac{1}{2d}G(x,x)\leq G^{\mathscr{E}(\omega)}(x,x)\leq\frac{1}{c(\mathscr{E},x)^{d}}G(x,x).$
This shows the conclusions of Pölya’s Theorem, which states
Theorem 3.6 (Pölya).
The simple symmetric Random Walk on $\mathbb{Z}^{d}$ is transient in $d\geq 3$ and recurrent in $d=1,2$.
carries on to any random walk in a balanced, elliptic environment, therefore answering Conjecture 1.2 in the affirmative. Recall that a random walk is transient or recurrent depending on $G^{\mathscr{E}({\omega})}(x,x)<\infty$ or $G^{\mathscr{E}({\omega})}(x,x)=\infty$.
@@ 3.1. Embedded Environments
Let $\mu$ be a stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.
@@
$\Lambda_{\omega}(x,v)$ $=\frac{\omega(x,v)+\omega(x,-v)}{2}$ (1)
@@
Assume $d\geq 1$. Let $\mu$ be an stationary ergodic law on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$. Then suppose the environment process for the Embedded Environment satisfies the conclusions of 3.1.
Then there exists a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic.
@@
Let $d\geq 1$, and let $\mu$ a stationary ergodic environment law on $\mathcal{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
@@ 3.2. Remarks:
The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$, namely the original and a reflected walk. This is motivated by the fact that the averaged random walk is balanced. Through the control argument, we pass from the balanced to the general case. The Action of Linear Transformations proves crucial in performing this step. This might and should appear strange without prior context, but we shall see this arising naturally as an artifact of our proof technique.
@@ 4. An outline of the Proofs
Let us begin this section by seeing why an invariant measure for the local process guarantees answers to Conjectures 1.1. We assume for now that Theorems 3.1 and 3.7 to be true and prove 3.8
@@
First note that by Theorems 3.1 and 3.7 we have a stationary ergodic measure on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic since Assumptions 2.2 and 2.3 are satisfied. Call this measure $\mathbb{Q}$. Assume without loss of generality that $X_{\omega}(0)=0$.
@@
And by Birkhoff’s Ergodic Theorem [2],
$\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to\int_{\mathcal{S}(d)}d(0,{\omega})d\mathbb{Q}\hskip 5.0pt\ \ \mathbb{Q}\ \ \text{a.e.}$
Combining both gives us the theorem $\mathbb{Q}_{0}=\mathbb{Q}\rtimes\mathbb{P}_{0}^{\omega}$ a.e., which by mutual absolute continuity of measures shows it holds $\mu^{0}=\mu\rtimes\mathbb{P}_{0}^{\omega}$ a.e.
@@
$(\Sigma_{k})_{ij}=\begin{cases}T(k)_{i}+T(k)_{i+d}-(T_{i}(k)-T_{i+d}(k))^{2}&\text{if }i=j,\\ -[T_{i}(k)-T_{i+d}(k)][T_{j}(k)-T_{j+d}(k)],&\text{if }i\neq j\end{cases}$
@@
The proof of Theorem 3.5 is postponed and will be shown as a natural consequence of our uniqueness principle. Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.7. We divide the proof into various sections in order to properly motivate the proof.
4.2. Discussion on Theorems 3.1, Corollary 3.2 and Theorem 3.5:
This section provides an alternate proof of Theorem 3.3 but generalises it to the general case of stationary, ergodic environments and arbitrary environment functions. This theorem is important for verifying the conditions of Theorem 3.7 under the assumptions of Theorem 3.8.
The two main themes in this proof are the use of Periodised Environments as used in [7] and our uniqueness principle. We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof. The proof in [7] adopts the arguments of Papanicolaou and Varadhan [8] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments. Our main step is to refine the arguments in [7] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle. The ideas of Krylov [6] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler. We adopt these ideas to bound a certain resolvent as in [7] under assumptions 2.2 and 2.3. We further use this uniqueness principle to derive Theorem 3.5. We emphasise that Theorem 3.5 is not a result on Random Walk in Random Environments but on general balanced elliptic walks.
@@
There are two core themes in this proof. The first being the use of Action of Linear Transformations on the Environment. This approach is novel to the best of my knowledge. Let us describe this theory in brief. When we talk of shifts on the environment $\tau_{x}\omega$, we recall it is given by the relation
@@
Through the Action theory, we talk of how general linear transformations on $\mathbb{Z}^{d}$ behave on the environment. Precisely speaking, we study expressions of the form $\tau_{Tx}\omega$ or $\tau_{x}T.\omega$, where $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$. These notions prove extremely fundamental to our arguments as we shall see.
We shall prepare the reader accordingly in the next section for the proof.
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. This technique allows us to control the paths of the random walk, which shall be key to the proof. The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. The construction is specific to the proof and shall be demonstrated accordingly during the proof.
The main hurdle in the proof is to show stationarity of the local process. The ergodicity of the local process follows by adapting arguments from [3] suitably, which shall be demonstrated.
5. Preparation for the Proof of Theorem 3.1, Corollary 3.2 and Theorem 3.5 :
Fix an environment ${\omega}\in\mathcal{S}(d)$ and $\mathscr{E}$ a balanced measurable environment function on $(\mathcal{S}(d),\mathcal{B}(d)^{\prime})$.

### v4

# Ergodic Theorems for Random Walks in Random Environments

@@
Abstract.
We study the ergodic properties of random walks in stationary ergodic environments without uniform ellipticity under a minimal assumption. There are two main components in our work. The first step is to adopt the arguments of Lawler, utilizing a more general definition of environments via environment functions, to deduce an invariance principle for balanced environments under some natural assumptions. Secondly, we transfer these results to general ergodic environments using a control technique to derive a measure under which the local process is stationary and ergodic. Two important consequences of our results are the Law of Large Numbers and the Invariance Principle.
keywords:
Random walk, Random environment, Ergodic Theorems
@@ 1. Introduction
The purpose of the discussion is to derive Ergodic Theorems and related results in the context of Random Walk in Random Environments (RWRE). This shall help us to mainly solve the problem of Random Walk in Random Environments regarding the Law of Large Numbers.
Roughly speaking, the first one says that there is a deterministic limiting velocity of the walk. This has been studied in detail in $d=1$ by Solomon[solomon] in 1975. In higher dimensions, specifically in $d\geq 3$, there is a lack of knowledge on both the existence and the explicit form of a Law of Large Numbers, which we shall answer in our setting.
Informally speaking, the problem regarding the Law of Large Numbers in RWRE, for stationary ergodic environments, is the following:
@@
$\frac{X_{n}}{n}\to v$
for almost every realisation of the environment?
Varadhan in [v] showed that the sequence $X_{n}/n$ has at most two deterministic limit points $v_{1},v_{2}$ and if $v_{1}\neq v_{2}$ then there is a constant $a\geq 0$ such that $v_{1}=-av_{2}$. Furthermore, Noam Berger in [ber] showed that for $d\geq 5$ if $v_{1}\neq v_{2}$ at least one of $v_{1},v_{2}$ is 0. These results assume that the environment is uniformly elliptic and i.i.d. In our present discussion, we answer 1.1 in the affirmative for stationary ergodic nearest neighbour RWRE and show the existence of the limiting velocity under some assumptions.
The 0-1 conjecture (which is a famous one in RWRE theory), first stated (for two dimensions) by Stephen Kalikow in [kalikow1981generalized] for uniformly elliptic i.i.d or more generally for uniformly elliptic ergodic environments, says that given a direction, the walk either escapes off or does not for almost all realisations of the environment. More formally
@@
Let $\ell\in S^{d-1}$. Let $A_{\ell}=\{X_{n}.\ell\to+\infty\}$. Then
$\mathbb{P}[A_{\ell}]\in\{0,1\}$
Prior work on this problem, notably in two dimensions, by Martin Zerner and Franz Merkl [merkl2001zero], has solved this for i.i.d. environments and has counterexamples for elliptic, ergodic environments. Further counterexamples exist in $d\geq 3$ in [bramson2006shortest] for stationary environments under a polynomial mixing condition, which is not equivalent to ergodic environments (it is not difficult to find counterexamples). The proofs crucially rely on invalidating the law of large numbers, so we expect, in view of our results, that this conjecture is also affirmative under our assumptions, but it should require a rigorous proof, which we are unable to provide.
We shall work under a more general definition of environments using Environment Function, which we shall introduce below. We shall establish under some minimal assumption that this chain has a stationary measure for balanced environments, adapting arguments from [Lawler1982] for any environment function. This shall also provide an easier and more natural proof of Theorem 1 of [guo2012quenched]. This shall be derived by adapting the arguments from [Lawler1982].
Our other main contribution will be to deduce the ergodicity of a certain stochastic process, which we will call the local process. This process is the projection at the origin of the “Environment Viewed from the Particle” Markov Chain, or, as we will call it in this paper, the environment process with a suitable environment function. We consider this process since it is sufficient for our purposes of deriving the law of large numbers and also in view of the fact that the environment process does not always possess suitable ergodic properties. This can be noted for Random Walks in Dirichlet Environments as studied by Christopher Sabot in [11].
Our definition of environments allows us to transfer our results from the balanced case to the general environment by coupling the random walk and a reflected walk. We then combine a novel technique, which we shall name as the Action of Linear Transformations, and another control technique to execute a proof of the law of large numbers.
As a consequence of our ergodic theorem, we shall also be able to deduce an invariance principle for general environments.
Our results hold for all uniformly elliptic RWRE in dimensions equal to or greater than one, and the conditions we impose show the extent to which the condition of uniform ellipticity can be dropped.
@@ 2. Prerequisites
Before stating the main results, we formally define the setting we work under. We adopt the formalism as in [Lawler1982].
@@
$\mathscr{S}(d)$ $=\{\text{set of functions from }\mathbb{Z}^{d}\to S(d)\}.$
Let $\mathcal{B}(d)$, $\mathcal{B}(d)^{\prime}$ be the Borel-$\sigma$ algebra on $S(d)$, $\mathscr{S}(d)$ respectively.
Consider a measurable function $\mathscr{E}:S(d)\to S(d)$. Call the environment function, the natural extension of this function on $\mathscr{S}(d)$ given by
$\mathscr{E}(\omega)(x)=\mathscr{E}(\omega(x)).$
We shall call such functions the environment function.
This function essentially contains the information on the transition probabilities of the random walk based on the function $\omega$. This is essentially a mild but powerful generalisation of the environment, as we shall see.
@@
Given an environment function $\mathscr{E}$, $\omega\in\mathscr{S}(d)$, we call the pair $(\omega,\mathscr{E})$ as the environment.
Let $\mu$ be a probability measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$. We call the environment function balanced if
@@
Define the action of the group of shifts $\{\tau_{x}\}_{x\in\mathbb{Z}^{d}}$ on $\omega\in\mathscr{S}(d)$ given by
@@
For $\mathscr{E}(\omega)\in\mathscr{S}(d)$, $v\in V_{d}$, $\mathscr{E}$ an environment function
@@
Fix $\omega\in\mathscr{S}(d)$ and an environment function $\mathscr{E}$. Define the Markov Chain of the Random Walk $\{X_{\mathscr{E}(\omega)}(k)\}_{k\geq 0}$ in $\mathbb{Z}^{d}$ by the transition probabilities,
@@
Define the environment process as the Markov Chain given by $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ with the transition kernel given by
@@
Therefore if ${\mathscr{E}(\omega)}\in\mathscr{S}(d)$ one has that
$\mathbb{P}^{{\mathscr{E}(\omega)}}_{x}(G):\mathscr{S}(d)\to[0,1]$
is measurable $\forall G\in\mathcal{G}$. Hence, we can define the annealed law to be the probability measure on $(\mathscr{S}(d)\times(\mathbb{Z}^{d})^{\mathbb{N}},\mathcal{B}(d)^{\prime}\times\mathcal{G)}$ as
@@
We make the following assumptions about the environment that is
@@
Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$. Since we aim to drop uniform ellipticity, i.e., all transition probabilities bounded below by some constant positive value, we define the quantity $c(\mathscr{E},x)$(we may drop $\omega$ from the notation) as follows
@@
We use the notion of concave functions defined on the compact $L_{1}$ ball $B_{1}[0,r]\cap\mathbb{Z}^{d}$. A function $z:B_{1}[0,r]\cap\mathbb{Z}^{d}\to\mathbb{R}_{\geq 0}$ $(r>0)$ is said to be concave if
@@
We also define an operator on concave functions, which we shall call the Monge-Ampère Operator. For a concave function $z$ we define the operator $M$ as
$Mz(x)=\prod_{i=1}^{d}[z(x+e_{i})+z(x-e_{i})-2z(x)].$
@@ 3. Main Results and Discussions
We first state a result on balanced environments. This result forms the backbone for our control argument as we shall see.
@@
Let $d\geq 1$. Assume $\mu$ is a stationary, ergodic Law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ with $\mathscr{E}$ a measurable balanced environment function satisfying assumptions 2.2 and 2.3.
Then there exists a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ such that the environment process $\{\tau_{X_{\mathscr{E}(\omega)}(k)}\omega\}_{k\geq 0}$ is stationary, ergodic under this measure and is mutually absolutely continuous to $\mu$.
@@
Let $d\geq 1$. Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ and $\mathscr{E}$ a measurable balanced environment function. Fix $\omega\in\mathscr{S}(d)$. Then under assumptions 2.2 and 2.3,
@@
These results notably weaken the assumptions of ellipticity and derive an alternate proof of Theorem 1 of [guo2012quenched] for $d\geq 1$, which states that
Theorem 3.3 (Theorem 1(i) of [guo2012quenched]).
@@
$\varepsilon_{\omega}(x)=\prod_{i=1}^{d}[\omega(x,e_{i})]^{\frac{1}{d}}.$
@@
To prove the above results, we shall need to bound a certain resolvent using martingale arguments and bounds on the Monge-Ampère Operator using arguments from 3.1.
@@ 3.1. Embedded Environments
Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.
@@
(1) $\Lambda_{\omega}(x,v)$ $=\frac{\omega(x,v)+\omega(x,-v)}{2}$
@@
The embedding function basically averages the probabilities of the opposing directions when the environment function is the Identity function.
@@
Assume $d\geq 1$. Let $\mu$ be a stationary ergodic law on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$. Then suppose the environment process for the Embedded Environment satisfies the conclusions of 3.1.
Then there exists a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic.
@@
Let $d\geq 1$, and let $\mu$ be a stationary ergodic environment law on $\mathscr{S}(d)$ with $\Lambda$ satisfying Assumption 2.2 and 2.3. Then:
@@ 3.2. Remarks:
The embedding function allows us to couple two random walks in $\mathbb{Z}^{d}$, namely the original and a reflected walk. This is motivated by the fact that the averaged random walk is balanced. Through the control argument, we pass from the balanced to the general case. The Action of Linear Transformations proves crucial in performing this step.
@@ 4. An outline of the Proofs
Assume without loss of generality that $X_{\mathscr{E}(\omega)}(0)=0$ where $\mathscr{E}$ is an environment function, for the proofs unless otherwise indicated.
Let us begin this section by seeing why an invariant measure for the local process guarantees answers to Conjectures 1.1. We assume for now that Theorems 3.1 and 3.4 are true and prove 3.5.
@@
First note that by Theorems 3.1 and 3.4 we have a stationary ergodic measure on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$ mutually absolutely continuous to $\mu$, such that the local process $\{\pi_{0}(\tau_{X_{\omega}(m)}\omega)\}_{m\geq 0}$ is stationary and ergodic since Assumptions 2.2 and 2.3 are satisfied. Call this measure $\mathbb{Q}$.
@@
And by Birkhoff’s Ergodic Theorem [birkhoff1931proof],
$\sum_{k=0}^{n-1}\frac{d(0,\tau_{X_{\omega}(k)}\omega)}{n}\to\int_{\mathscr{S}(d)}d(0,{\omega})d\mathbb{Q}\hskip 5.0pt\ \ \mathbb{Q}\ \ \text{a.e.}$
Combining both gives us the theorem $\mathbb{Q}^{0}=\mathbb{Q}\rtimes\mathbb{P}_{0}^{\omega}$ a.e., which by mutual absolute continuity of measures shows that it holds $\mu^{0}=\mu\rtimes\mathbb{P}_{0}^{\omega}$ a.e.
@@
$(\Sigma_{k})_{ij}=\begin{cases}T_{i}(k)+T_{i+d}(k)-(T_{i}(k)-T_{i+d}(k))^{2}&\text{if }i=j,\\ -[T_{i}(k)-T_{i+d}(k)][T_{j}(k)-T_{j+d}(k)],&\text{if }i\neq j\end{cases}$
@@
Thus, we shall devote the rest of the paper to proving Theorems 3.1 and Theorem 3.4. We divide the proof into various sections in order to properly motivate the proof.
4.2. Discussion on Theorems 3.1, Corollary 3.2:
This section provides an easier proof of Theorem 3.3 but generalises it to the general case of stationary, ergodic environments and arbitrary environment functions. This theorem is important for verifying the conditions of Theorem 3.4 under the assumptions of Theorem 3.5.
The two main themes in this proof are the use of Periodised Environments as used in [Lawler1982] and a uniqueness principle for concave functions. We shall prepare the reader with the prerequisites for the proof in the next section and fully motivate the proof. The proof in [Lawler1982] adopts the arguments of Papanicolaou and Varadhan [PapanicolaouVaradhan1982] for diffusion processes with random coefficients for the random walk for uniformly elliptic balanced environments. Our main step is to refine the arguments in [Lawler1982] for non-uniformly elliptic balanced environments under assumptions 2.2 and 2.3 to get our uniqueness principle. The ideas of Krylov [Krylov1971] were used to estimate the solutions of the discrete Monge-Ampère Equation by Lawler. We adopt these ideas to bound a certain resolvent as in [Lawler1982] under assumptions 2.2 and 2.3.
@@
There are two core themes in this proof. The first being the use of Action of Linear Transformations on the Environment. This approach is novel in the sense that it generalises the affine shift operator on the environment to linear transforms. Let us describe this theory in brief. When we talk of shifts on the environment $\tau_{x}\omega$, we recall it is given by the relation
@@
Through the Action theory, we talk of how general linear transformations on $\mathbb{Z}^{d}$ behave on the environment. Precisely speaking, we study expressions of the form $\tau_{Tx}\omega$ or $\tau_{x}T.\omega$, where $T\in\mathrm{Hom}_{\mathbb{Z}}(\mathbb{Z}^{d},\mathbb{Z}^{d})$. These notions prove extremely fundamental to our arguments as we shall see. We shall prepare the reader accordingly in the upcoming sections for the proof.
The second core theme in this proof is to pass from the embedded environment to the original environment by exhibiting stationarity of the local process through a control argument using Bernoulli Random Variables Independent of the environment. This allows us to “couple” the original walk and a reflected walk. This technique allows us to control the paths of the random walk, which shall be key to the proof. The control argument will therefore allow us to use the stationary measure in $\mathbb{Z}^{d}$ in the Embedded Environment and pass on to the environment in $\mathbb{Z}^{d}$ to show stationarity of the local process. In particular, the local process of the walk shall be equal in law to the local process of the random walk in the embedded environment. The control argument shall help us achieve this by sort of averaging the original and the reflected, as might be noted in the choice of the embedding function.
The main hurdle in the proof is to show stationarity of the local process. The ergodicity of the local process follows by adapting arguments from [BolthausenSznitman2002] suitably, which shall be demonstrated.
5. Preparation for the Proof of Theorem 3.1, Corollary 3.2 :
Fix an environment ${\omega}\in\mathscr{S}(d)$ and $\mathscr{E}$ a balanced measurable environment function on $(\mathscr{S}(d),\mathcal{B}(d)^{\prime})$.